Problem 11-1A Short-term notes payable transactions and entries LO P1

[The following information applies to the questions displayed below.]

Tyrell Co. entered into the following transactions involving short-term liabilities in 2016 and 2017.

2016

Apr. 20 Purchased $35,500 of merchandise on credit from Locust, terms n/30. Tyrell uses the perpetual inventory system.
May 19 Replaced the April 20 account payable to Locust with a 90-day, $35,000 note bearing 7% annual interest along with paying $500 in cash.
July 8 Borrowed $63,000 cash from NBR Bank by signing a 120-day, 11% interest-bearing note with a face value of $63,000.
__?__ Paid the amount due on the note to Locust at the maturity date.
__?__ Paid the amount due on the note to NBR Bank at the maturity date.
Nov. 28 Borrowed $24,000 cash from Fargo Bank by signing a 60-day, 6% interest-bearing note with a face value of $24,000.
Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank.


2017

__?__ Paid the amount due on the note to Fargo Bank at the maturity date.

Problem 11-1A Part 1

Required:
1. Determine the maturity date for each of the three notes described.

Locust NBR Bank Fargo Bank Maturity date

Problem 11-1A Part 2

2. Determine the interest due at maturity for each of the three notes. (Do not round your intermediate calculations. Use 360 days a year.)

Principal X Rate x Time - Interest Locust NBR Bank Fargo Bank

Problem 11-1A Part 3

3. Determine the interest expense to be recorded in the adjusting entry at the end of 2016. (Do not round your intermediate calculations. Use 360 days a year.)

Year end accrual required for: Fargo Bank Principal X Rate x Time = Interest Interest to be accrued in 2016
Answer #1

1) Maturity date

Locust NBR Bank Fargo Bank
Maturity date 17 August 5 November 27 January

2) calculate interest

Principal * Rate * Time = Interest
Locust 35000 * 7% * 90/360 = 613
NBR Bank 63000 * 11% * 120/360 = 2310
Fargo Bank 24000 * 6% * 60/360 = 240

3) Year end accrual required for

Principal * rate * Time = Interest
Interest to be accrued in 2016 24000 * 6% * 33/360 = 132